# Integrating IP broadcasting with audio tags: workflow and Challenges

### v2

@@ 6 Discussion and Conclusion
One of the primary challenges discussed is the latency associated with the audio tagging model. Given the real-time nature of broadcasting, any delays introduced by processing can impact the overall operation. This makes the choice of buffer size crucial. A smaller buffer reduces latency but might compromise the accuracy of sound event detection. Conversely, a larger buffer improves accuracy but increases latency. The experiments conducted show that an acceptable balance can be achieved with a buffer size of 48128 samples, which provides an acceptable latency while maintaining accuracy. The use of Efficient PANNs (E-PANNs) further helps in reducing the computational complexity and memory requirements, making it a suitable choice for real-time applications.
Containerisation offers a robust solution to scalability issues. By isolating the audio tagging functionality into a microservice, it becomes possible to scale the system by simply adding more containers as needed. This isolation also ensures that a fault in one container does not affect the entire system, enhancing overall reliability. The use of Docker to containerise these services allows for easy deployment and management across different network setups. Additionally, the integration with NDI technology, which is widely adopted in the industry, ensures broad applicability.
@@ 6.1 Conclusion
Integrating IP broadcasting with audio tagging presents a promising advancement for the broadcasting industry. The use of containerisation and audio tagging for real-time sound event detection can significantly enhance content production and accessibility. However, addressing the challenges of latency, accuracy and real-world deployment is crucial for the successful implementation of this technology. Future work includes re-writing the codebase to use the NDI C++ SDK directly, avoiding the issues surrounding the Python bindings. Additionally, we would like to analyse more complex models such as transformers [17, 18] within our broadcasting framework. Finally, the creation of the discussed proof of concept applications would allow for full demonstration of the usefulness of this technology.
@@ 7 Acknowledgements
This work was supported by Engineering and Physical Sciences Research Council (EPSRC) Grant EP/T019751/1 “AI for Sound (AI4S)”. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.

### v3

@@ 6 Discussion and Conclusion
One of the primary challenges discussed is the latency associated with the audio tagging model. Given the real-time nature of broadcasting, any delays introduced by processing can impact the overall operation. This makes the choice of buffer size crucial. A smaller buffer reduces latency but might compromise the accuracy of sound event detection. Conversely, a larger buffer improves accuracy but increases latency. Experiments conducted (Table 1) show that an acceptable balance can be achieved with a buffer size of 48128 samples, which provides an acceptable latency while maintaining accuracy. The use of Efficient PANNs (E-PANNs) further helps in reducing the computational complexity and memory requirements, making it a suitable choice for real-time applications.
Containerisation offers a robust solution to scalability issues. By isolating the audio tagging functionality into a microservice, it becomes possible to scale the system by simply adding more containers as needed. This isolation also ensures that a fault in one container does not affect the entire system, enhancing overall reliability. The use of Docker to containerise these services allows for easy deployment and management across different network setups. Additionally, the integration with NDI, which is widely adopted in the industry, ensures broad applicability.
@@ 6.1 Conclusion
Integrating IP broadcasting with audio tagging presents a promising advancement for the broadcasting industry. The use of containerisation and audio tagging for real-time sound event detection can significantly enhance content production and accessibility. However, addressing the challenges of latency, accuracy and real-world deployment is crucial for the successful implementation of this approach. Future work includes re-writing the codebase to use the NDI C++ SDK directly, avoiding the issues surrounding the community Python bindings. Additionally, we would like to analyse more complex models such as transformers [17, 18] within our broadcasting framework. Finally, the creation of the discussed proof of concept applications would allow for full demonstration of the usefulness of this technology.
@@ 7 Acknowledgements
This work was supported by the Engineering and Physical Sciences Research Council [grant number EP/T019751/1] and an Adobe Research Gift. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.